What is the value of x if x/5=x/6+1
Answer by algebrahouse.com(1659)

"What is the value of x if x/5=x/6+1"

x/5 = x/6 + 1
30[x/5 = x/6 + 1) {multiplied entire equation by least common denominator, 30}
30(x/5) = 30(x/6) + 30(1) {multiplied each term by 30}
6x = 5x + 30 {multiplied through}
x = 30 {subtracted 5x from both sides}
www.algebrahouse.com